a)      Canadian Bacon Inc. financial statements are presented in the table below.

       Based on the information in the table, and using cost of goods sold and a 365-day year, calculate Days of Sales in Inventory (using cost of goods sold).

Round the answers to two decimal places

Balance Sheet December 31, 2012

Cash and marketable securities $198,000 Accounts payable $288,000
Accounts receivable $469,000 Notes payable $65,000
Inventories $577,000 Accrued expenses $84,000
Prepaid expenses $15,700 Total current liabilities $437,000
Total current assets $1,259,700 Long-term debt $237,000
Gross fixed assets $1,954,000 Par value and paid-in-capital $199,000
Less: accumulated depreciation $476,000 Retained Earnings $1,864,700
Net fixed assets $1,478,000 Common Equity 2,063,700
Total assets $2,737,700 Total liabilities and owner’s equity $2,737,700

Income Statement, Year of 2012

Net sales (all credit) $7,546,600.00
Less: Cost of goods sold $6,112,746.00
Selling and administrative expenses $349,000.00
Depreciation expense $145,000.00
EBIT $939,854.00
Interest expense $49,500.00
Earnings before taxes $890,354.00
Income taxes $356,141.60
Net income $534,212.40
Answer #1

Cost of goods sold is $6,112,746.00 from the income statement.

The question states us to take 365 days per year,

Average day's COGS = Cost of goods sold / 365 days = 6,112,746/365 = 16747.2493151

inventories from balance sheet are $577,000

Thus, days of sale in inventories is Inventories / average day's COGS

=577000/16747.2493151

=34.45 days